OpenStudy (anonymous):

(PLEASE HELP) Suppose the base and height of a triangle to 20, and the area of the triangle is 42 square cm. What are the dimensions (base and height) of the triangle? Need help starting off well enough to get the answer. so far im stuck on 42=1/2b(20-b)

OpenStudy (wolf1728):

You need to set up 2 equations A) (height * base)/2 = 42 B) height + base = 20 rearranging equation B base = 20 - height Substituting this into equation A [height * (20 -height)]/2 = 42 -height^2 + 20 * height = 42 -height^2 + 20 * height -42 = 0 using quadratic formula where a = -1 b = +20 c=-42 height = 2.3842 and height = 17.616 therefore base = 2.3842 and 17.616

OpenStudy (anonymous):

wow thank you I wrote that down for later reference thanks!

OpenStudy (wolf1728):

u r welcome :-)